Following is partial information for Delamunte Industries for the month of August:

Work In Process
Balance, August 1 $ 134,000
Direct materials used 114,800
Direct labor 163,000
Manufacturing overhead applied
(based on direct labor cost)
152,000

Jobs finished during August are summarized here:

Job # Cost of Jobs Completed
234 $ 59,950
237 28,800
231 47,000
246 20,000

Required:
At the end of August, only one job, Job 248, was still in process. The direct labor cost incurred on Job 248 as of August 31 was $24,300.


1. Calculate the predetermined overhead rate that was used to apply manufacturing overhead to jobs during August.

2. Determine the balance in Work in Process at the end of August.

3. Determine how much of the direct materials cost and applied overhead would be recorded for Job 248 as of August 31
Answer #1
Answer

1

Predetermined overhead rate = 152000/163000= $0.93 per Direct labor Dollar
2
Total manufacturing costs = 134000+114800+163000+152000= $563,800
Total cost of finished jobs = 59950+28800+47000+20000= $155,750
Work in Process at the end of August = 563,800- 155,750 = $408050
3
Applied overhead = 24,300*0.93= $22,599
Direct materials = 408050-24,300-22,599= $361,151